A 6.35 L sample of carbon dioxide is collected at 550 ° C and 90.312 kPa. What volume will the gas occupy at 106.391 kPa and 20.0 ° C?


1
Expert's answer
2022-01-13T03:55:02-0500

"\frac{P_1×V_1}{T_1}=\frac{P_2×V_2}{T_2}=nR\\V_2=\frac{P_1×V_1×T_2}{T_1×P_2}=\frac{n×R×T_2}{P_2}\\=\frac{90.312×6.35×293}{773×106.391}=2.043L
\\
\\"